# Vecchia Gaussian Processes: Probabilistic Properties, Minimax Rates and Methodological Developments

### v3

@@ B.4 Proof of Lemma 6
$\|K_{\nu A,\nu A}-K_{\nu_{0}A,\nu_{0}A}\|_{1}\lesssim|\nu-\nu_{0}|^{2\alpha},\quad\|K_{\nu A,\nu x}-K_{\nu_{0}A,\nu_{0}x}\|_{1}\lesssim|\nu-\nu_{0}|^{2\alpha}.$ (57)
@@
$\|K_{\nu A,\nu A}^{-1}-K_{\nu_{0}A,\nu_{0}A}^{-1}\|_{1}\leq\|K_{\nu_{0}A,\nu_{0}A}^{-1}\|_{1}\|K_{\nu A,\nu A}-K_{\nu_{0}A,\nu_{0}A}\|_{1}\|K_{\nu A,\nu A}^{-1}\|_{1}\lesssim|\nu-\nu_{0}|^{2\alpha}.$ (58)
Combining the inequalities (57) and (58), results in
@@
providing (32) in case $\alpha\leq 1$.
@@
Proof of (31): Let us write $M_{\nu}=O(\nu^{\delta})$ for a matrix $M_{\nu}$ indexed by a parameter $\nu$ and $\delta\geq 0$ if $\|M_{\nu}\|_{1}\lesssim\nu^{\delta}$ as $\tau$ tends to 0. We recall that for finite dimensional matrices, the matrix $\ell^{1}$ norm is equivalent to various other matrix norms, including $\ell^{2}$, $\ell^{\infty}$ and trace norms.
@@
$B_{j,x}:=V_{A}^{-1}V_{A_{j,x}}.$ (60)
Note that except of the $j$th column, the matrix $B_{j,x}$ coincides with the identity matrix, while its $j$th column is $V_{A}^{-1}v_{x}$. Therefore,
@@
$\frac{\det(K_{\nu A,\nu A_{j,x}})}{\det(K_{\nu A,\nu A})}=\det(B_{j,x})+O(\nu+\nu^{2(\alpha-\underline{\alpha})}).$ (61)
The combination of the previous four displays imply that
@@
Hence it remained to show that (61) holds. We proceed by reformulating the matrices $K_{\tau A,\tau A_{j,x}}$ and $K_{\tau A,\tau A}$ into a more suitable, product form, by taking their Taylor expansions. We also introduce some additional notations to facilitate the arguments.
@@
$K(x_{1},x_{2})=\sum_{|k_{1}|+|k_{2}|\leq 2\underline{\alpha}}q_{k_{1},k_{2}}x_{1}^{k_{1}}x_{2}^{k_{2}}+O\big{(}(\|x_{1}\|_{2}+\|x_{2}\|_{2})^{2\alpha}\big{)},$ (62)
@@
Then taking the Taylor expansion (62) for each element of the matrix $K_{\tau A,\tau A}$ yields
@@
$=$ $V_{A}^{T}\Gamma_{1}\tilde{Q}_{11}\Gamma_{1}V_{A}\big{[}I+O(\nu^{2(\alpha-\underline{\alpha})})\big{]},$ (63)
@@
Next we proceed with reformulating the covariance matrix $K_{\tau A,\tau A_{j,x}}$ in a similar manner. First note, that in view of Lemma 2 and its proof in Section A.2, for a norming set $A$ with norming constant $c_{N}$, for all $P\in\mathscr{P}_{l}([0,1]^{d})$,
@@
where $R_{j,x}=\tilde{V}_{A_{j,x}}-\tilde{V}_{A}B_{j,x}$. Since $A\cup\{x\}\subset[0,1]^{d}$, in view of (64), the error term $R_{j,x}$ has bounded $\ell^{1}$ norm independent of the set $A$ and point $x$. Then, the covariance matrix $K_{\tau A,\tau A_{j,x}}$ can be written as
@@
$=$ $V_{A}^{T}\Gamma_{1}\tilde{Q}_{11}\Gamma_{1}V_{A}\big{[}B_{j,x}+O(\tau)+O(\tau^{2(\alpha-\underline{\alpha})})\big{]}.$ (65)
Since the determinat is a multiplicative map, we get in view of the previous display and (B.4) that
@@
concluding the proof of (61).
Finally, in case the origin doesn’t belong to the convex hull of $A\cup\{x\}$, then take an arbitrary $x_{0}\in A\cup\{x\}$, and note that $0\in A\cup\{x\}-x_{0}$. Hence, by working with the shifted process $(Z_{x-x_{0}}^{\tau,s}|Z_{A-x_{0}}^{\tau,s})$ the conclusion holds. ∎
B.5 Bounding $\vartheta_{n}$ defined in (34)
The quantity $\vartheta_{n}$ plays a crucial role in controlling the recursive polynomial and Gaussian interpolations and hence in deriving small deviation bounds for Vecchia GPs. We focus on grid data and verify that for $d\leq 2$ and not too large regularity parameter $\alpha$, $\vartheta_{n}=O(1)$. Controlling this term in higher dimensions or more general design settings is highly challenging and going beyond the scope of the present paper. But we conjecture that the results hold more generally than our restricted setting. To simplify the analysis, we consider a slightly modified version of Algorithm 3, to deal with the data points close to the boundary. We discuss this technical extension in the next subsection in details.
Lemma 14.
Suppose Condition 2 holds, $\mathcal{X}_{n}$ is on a grid and the DAG of the Vecchia GP is constructed by Algorithm 3. If one of the following holds:
@@ B.5 Bounding $\vartheta_{n}$ defined in (34)
We start with discussing the conditions of Lemma 14. Although the range of $d$ and $\alpha$ may look somewhat restricted at first glance, they in fact cover the cases typically considered in geostatistical applications, where Vecchia GP is routinely used. In such applications, the domain is a geological space with dimension $d$ equals $1$ or $2$ and the regularity parameter $\alpha$ typically does not exceed $5/2$.
The main idea in the proof follows from the observation that in case of infinite grid, the geometric shapes of $\mathrm{pa}(X_{i})$ are the same. Therefore, the polynomial interpolation operators $P_{j},\forall j$ can be viewed as convolution operators, which can be further converted to pointwise multiplication via Fourier transform. This approach will greatly simplify our analysis. However, we consider a finite grid which at the boundary has different behaviour than the infinite grid. Polynomial interpolation on a finite grid close to the boundary has “boundary effects”, which has to be dealt with separately. To this end we will use data augmentation. We expand the dataset $\mathcal{X}_{n}$ with artificially placed locations, such that any $X_{i}\in\mathcal{X}_{n}$ and its ancestors are sufficiently far from the boundary. We proceed by first embedding the finite grid $\mathcal{X}_{n}$ into another, large enough grid and then construct a Layered Norming DAG on the expanded grid similarly to Algorithm 1. We make this general strategy precise in the rest of the section.
Our proof is divided into three parts. The first part discusses how to extend a gird to eliminate boundary effects. The second part defines the polynomial interpolation on an infinite grid and shows that it is equivalent to a finite polynomial interpolation on the extended grid. In the final part we provide upper bounds for the polynomial interpolation on infinite grid using Fourier analysis.
B.5.1 Extension of grid
We start by introducing some notations. Without loss of generality, let the dataset $\mathcal{X}_{n}$ be a $n=(2^{r}+1)^{d}$ grid for some $r\in\mathbb{N}$ and choose $\gamma=2$, i.e.
@@ B.5.1 Extension of grid
Observe, that the above DAG structure satisfies Condition 1 with $c_{d}=1$ and Condition 2 with a constant $c_{N}>0$. Next, we augment each $\mathcal{N}_{j}$ with elements from $\mathcal{G}_{j}$ to form new sets $\tilde{\mathcal{N}}_{j}$, such that none of the elements $x\in\mathcal{X}_{n}$ have boundary effects. Specifically, for all $X_{i}\in\mathcal{X}_{n}$, let $\mathcal{A}(X_{i})$ be the collection of all ancestors of $X_{i}$, i.e., the collection of all elements of $\mathcal{X}_{n}$ such that there exists a directed path from that element to $X_{i}$. Then for all $X\in\mathcal{A}(X_{i})\cap\mathcal{N}_{j}$, $j<\eta(i)$, under Condition 2 and recalling $\gamma=2$, we have
$\|X_{i}-X\|_{\infty}\leq\sum_{j^{\prime}=j}^{\eta(i)}c_{L}\gamma^{-j^{\prime}}\leq 2c_{L}\cdot 2^{-j}.$ (66)
Next we describe the iterative data augmentation procedure. Recall that in Algorithm 1 the parent sets for elements in layers $j\leq j_{0}$, with $j_{0}$ only depending on $\alpha$ and $d$, were initialized as $\mathrm{pa}(X_{i})=\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$. In this case we do not use data augmentation and set $\tilde{\mathcal{N}}_{j}=\mathcal{N}_{j}$, $\forall\;,0\leq j\leq j_{0}$. For all $j_{0}<j\leq\eta(n)-1$, we iteratively augment each layer $\mathcal{N}_{j}$ to $\tilde{\mathcal{N}}_{j}$ as
@@
is defined in the same way as in Algorithm 1 and satisfies Condition 2. This also means the relation (66) holds for all $X_{i}\in\tilde{\mathcal{X}}$ as well.
For all $X_{i}\in\tilde{\mathcal{X}}_{n}$ and $j_{0}\leq j\leq\eta(i)-1$, the new layer $\tilde{\mathcal{N}}_{j}$ includes all grid points in $\mathcal{G}_{j}$ within $4c_{L}\cdot 2^{-j}-4c_{L}\cdot 2^{-\eta(i)}\geq 2c_{L}\cdot 2^{-j}$ distance of $X_{i}$. In view of (66) all potential ancestors of $X_{i}$ in the $j$th layer is within $2c_{L}\cdot 2^{-j}$ distance of $X_{i}$, hence the DAG structure of Algorithm 3 eliminates all boundary effects for $X_{i}\in\cup_{j=j_{0}}^{\eta(n)}\tilde{\mathcal{N}}_{j}\subset\tilde{\mathcal{X}}_{n}$.
1 Input grid data $\mathcal{X}_{n}$ with $n=\tilde{n}^{d}$ and order $l$. Set $m={l+d\choose l}$ and let $r$ be the largest integer such that $\tilde{n}\geq 2^{r}+1$.
@@
10   Augment $\mathcal{N}_{j}$ to $\tilde{\mathcal{N}}_{j}$ according to (67).
@@
20         Order the projection of $\cup_{0\leq j^{\prime}\leq j-1}\tilde{\mathcal{N}}_{j^{\prime}}$ onto the $h$th coordinate by the increasing distance to $X[h]$ as $B_{h}=(x_{s(1),h},x_{s(2),h},\cdots,x_{s(n^{\prime}),h})$.
@@
We note that instead of data augmentation, one could also trim the data set to get rid of the boundary effect, losing only a negligible (tending to zero) fraction of the data. This can be done in an analogous way to data augmentation and is left as an exercise to the reader.
B.5.2 Polynomial interpolation on infinite grid
With the boundary effects eliminated by data augmentation, we can define the polynomial interpolation operator on infinite grids and align it with the $P_{j},j\in\mathbb{N}$, the operators on finite grids.
@@ B.5.2 Polynomial interpolation on infinite grid
$\mathbb{K}=\{k=(k[1],k[2],\cdots,k[d)]^{T}\in\mathbb{Z}^{d}:|k|\geq 1,0\leq k_{j}\leq 1\}.$ (68)
@@
Furthermore, let us denote by $\mathrm{pa}(k/2)$ the corner parent set defined in (21) for $k/2\in\mathcal{N}_{j}$ and $\cup_{0\leq j^{\prime}\leq j-1}\mathcal{N}_{j^{\prime}}$ a sufficiently large subset of $\mathbb{Z}^{d}$. Recalling the definition of the Vandermonde matrix $V_{A}$ in (14), we define a linear operator $L_{k}$ on functions $f:\mathbb{Z}^{d}\rightarrow\mathbb{R}$ as
$L_{k}(f)(z)=v_{k/2}^{T}V_{\mathrm{pa}(k/2)}^{-1}\bm{f}(z+\mathrm{pa}(k/2))\qquad\forall z\in\mathbb{Z}^{d},$ (69)
@@
$P(f)(z)=\Bigg{\{}\begin{aligned} &f(z/2),&&\text{if}\,z[j]\text{ is even},\;\forall 1\leq j\leq d,\\ &L_{k}(f)(\lfloor z/2\rfloor),\;&&\text{else, with $k=z-2\lfloor z/2\rfloor$, }\end{aligned}$ (70)
@@
This part aligns an arbitrarily function $f$ defined on the finite grid $\tilde{\mathcal{X}}_{n}$ with an appropriate function $f^{\prime}$ defined on an infinite grid $\mathbb{Z}^{d}$, such that under proper scaling, these two functions obtain the same polynomial interpolation over the grid.
Specifically, for all $0\leq j_{1}\leq j_{2}\leq\eta(n)-1$ and $f\in\ell^{1}(\tilde{\mathcal{N}}_{j_{1}})$, the objective of this paragraph is to find an integer $j_{3}\in\mathbb{N}$ satisfying $j_{1}\leq j_{3}\leq j_{2}$ and a function $f^{\prime}\in\ell^{1}(\mathbb{Z}^{d})$, such that
$P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)(x)=P^{j_{2}-j_{3}}(f^{\prime})(2^{j_{2}+1}x),\;\;\forall x\in\tilde{\mathcal{X}}_{n}\cap\mathcal{G}_{j_{2}+1},$ (71)
@@
$\|f^{\prime}\|_{\infty}\lesssim\|f\|_{\infty}.$ (72)
@@
Let $j_{3}=j_{1}$, then (72) is immediate. Furthermore, in view of Algorithm 3, for all $X_{i}\in\cup_{j=j_{1}}^{\eta(n)}\tilde{\mathcal{N}}_{j}\subset\cup_{j=j_{0}}^{\eta(n)}\tilde{\mathcal{N}}_{j}$ its ancestors are in $\tilde{\mathcal{X}}_{n}$. Therefore, $X_{i}$ has no boundary effect, implying directly (71).
For $j_{1}<j_{0}$, let us define an intermediate function $f_{j_{0}}\in\ell^{1}(\cup_{j=0}^{j_{0}}\tilde{\mathcal{N}}_{j})$ as
@@
Let $j_{3}=j_{0}$, then $\|f^{\prime}\|_{\infty}=\|f_{j_{0}}\|_{\infty}$ holds. Furthermore, since $j_{0}$ only depends on $d,\alpha$ and $P_{j},\forall j$ has bounded $\ell^{\infty}$ norm, we also have $\|f_{j_{0}}\|_{\infty}\lesssim\|f\|_{\infty}$. Thus, the upper bound (72) is satisfied. Furthermore, since for arbitrary $X_{i}\in\cup_{j=j_{0}}^{\eta(n)}\tilde{\mathcal{N}}_{j}$ there is no boundary effect, equation (71) also holds.
@@
Assertions (71) and (72) align a function defined on the finite grid $\tilde{\mathcal{X}}_{n}$ with an appropriate one defined on the infinite grid $\mathbb{Z}^{d}$. The polynomial interpolation operator $P$ has a localization property in the sense that the value of $P(f)(x)$ is only affected by $f(x^{\prime})$ for a limited number of $x^{\prime}$ that are close to $x$. Build on this idea, by introducing a surrogate function defined on $\mathbb{Z}^{d}$, we can derive a sharp bound for the operator norms that are suitable to study with Fourier analysis. We provide the details below.
For $j_{3}$ and $f^{\prime}\in\ell^{1}(\mathbb{Z}^{d})$ satisfying (71) and (72), and $X_{i}\in\tilde{\mathcal{X}}_{n}$, define the function $f_{i}^{\prime}\in\ell^{1}(\mathbb{Z}^{d})$ as
@@
In view of assertion (66) , the set $\{x:\|x-X_{i}\|_{\infty}\leq 2c_{L}\cdot 2^{-j_{3}}\}$ contains all the ancestors of $X_{i}$ that belong to the layers $\cup_{j=j_{3}}^{j_{2}-1}\tilde{\mathcal{N}}_{j}$.
@@
$P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)(X_{i})=P^{j_{2}-j_{3}}(f^{\prime})(2^{j_{2}+1}X_{i})=P^{j_{2}-j_{3}}(f_{i}^{\prime})(2^{j_{2}+1}X_{i}),$ (73)

### v4

@@ E.4 Proof of Lemma 3
$\|K_{\nu A,\nu A}-K_{\nu_{0}A,\nu_{0}A}\|_{1}\lesssim|\nu-\nu_{0}|^{2\alpha},\quad\|K_{\nu A,\nu x}-K_{\nu_{0}A,\nu_{0}x}\|_{1}\lesssim|\nu-\nu_{0}|^{2\alpha}.$ (E.11)
@@
$\|K_{\nu A,\nu A}^{-1}-K_{\nu_{0}A,\nu_{0}A}^{-1}\|_{1}\leq\|K_{\nu_{0}A,\nu_{0}A}^{-1}\|_{1}\|K_{\nu A,\nu A}-K_{\nu_{0}A,\nu_{0}A}\|_{1}\|K_{\nu A,\nu A}^{-1}\|_{1}\lesssim|\nu-\nu_{0}|^{2\alpha}.$ (E.12)
Combining the inequalities (E.11) and (E.12) results in
@@
providing (27) for the case $\alpha\leq 1$.
@@
Proof of (26): For $\delta\geq 0$, let us write $M_{\nu}=O(\nu^{\delta})$ for a matrix $M_{\nu}$ indexed by a parameter $\nu$, if $\|M_{\nu}\|_{1}\lesssim\nu^{\delta}$ as $\tau$ tends to 0. We recall that for finite dimensional matrices, the matrix $L_{1}$ norm is equivalent to various other matrix norms, including $_{2}, LL_{\infty}$ and trace norms.
@@
$B_{j,x}:=V_{A}^{-1}V_{A_{j,x}}.$ (E.14)
Note that except for the $j$th column, the matrix $B_{j,x}$ coincides with the identity matrix, while its $j$th column is $V_{A}^{-1}v_{x}$. Therefore,
@@
$\frac{\det(K_{\nu A,\nu A_{j,x}})}{\det(K_{\nu A,\nu A})}=\det(B_{j,x})+O(\nu+\nu^{2(\alpha-\underline{\alpha})}).$ (E.15)
The combination of the previous four displays implies that
@@
Hence, it only remains to show that (E.15) holds. We proceed by reformulating the matrices $K_{\tau A,\tau A_{j,x}}$ and $K_{\tau A,\tau A}$ into a more suitable, product form, by taking their Taylor expansions. We also introduce some additional notation to facilitate the arguments.
@@
$K(x_{1},x_{2})=\sum_{|k_{1}|+|k_{2}|\leq 2\underline{\alpha}}q_{k_{1},k_{2}}x_{1}^{k_{1}}x_{2}^{k_{2}}+O\big((\|x_{1}\|_{2}+\|x_{2}\|_{2})^{2\alpha}\big),$ (E.16)
@@
Taking the Taylor expansion (E.16) for each element of the matrix $K_{\tau A,\tau A}$ yields
@@
$=$ $V_{A}^{T}\Gamma_{1}\tilde{Q}_{11}\Gamma_{1}V_{A}\big[I+O(\nu^{2(\alpha-\underline{\alpha})})\big],$ (E.17)
@@
Next we proceed with reformulating the covariance matrix $K_{\tau A,\tau A_{j,x}}$ in a similar manner. First note, that in view of Lemma B.2 and its proof thereafter, for a norming set $A$ with norming constant $c_{N}$, for all $P\in\mathscr{P}_{l}([0,1]^{d})$,
@@
where $R_{j,x}=\tilde{V}_{A_{j,x}}-\tilde{V}_{A}B_{j,x}$. Since $A\cup\{x\}\subset[0,1]^{d}$, in view of (E.18), the error term $R_{j,x}$ has bounded $L_{1}$ norm independent of the set $A$ and point $x$. Then, the covariance matrix $K_{\tau A,\tau A_{j,x}}$ can be written as
@@
$=$ $V_{A}^{T}\Gamma_{1}\tilde{Q}_{11}\Gamma_{1}V_{A}\big[B_{j,x}+O(\tau)+O(\tau^{2(\alpha-\underline{\alpha})})\big].$ (E.19)
Since the determinant is a multiplicative map, we obtain, in view of the previous display and (E.4) that
@@
concluding the proof of (E.15).
Finally, in case the origin doesn’t belong to the convex hull of $A\cup\{x\}$, take an arbitrary $x_{0}\in A\cup\{x\}$, and note that $0\in A\cup\{x\}-x_{0}$. Hence, by working with the shifted process $(Z_{x-x_{0}}^{\tau,s}|Z_{A-x_{0}}^{\tau,s})$, the conclusion holds. ∎
E.5 Bounding $\vartheta_{n}$ defined in (29)
The quantity $\vartheta_{n}$ plays a crucial role in controlling the recursive polynomial and Gaussian interpolations, and hence in deriving small deviation bounds for Vecchia GPs. We focus on grid data and verify that, for $d\leq 2$ and moderate values of the regularity parameter $\alpha$, $\vartheta_{n}=O(1)$. Controlling this term in higher dimensions or for more general design settings is highly challenging and beyond the scope of the present paper. Nevertheless, we conjecture that the results extend beyond the restricted setting considered here. To simplify the analysis, we consider a slightly modified version of Algorithm 4 to handle data points near the boundary. We discuss this technical extension in detail the next subsection.
Lemma E.1.
Suppose Condition 2 holds, $\mathcal{X}_{n}$ is on a grid and the DAG of the Vecchia GP is constructed according to Algorithm 4. If one of the following conditions is satisfied:
@@ E.5 Bounding $\vartheta_{n}$ defined in (29)
We first discuss the conditions of Lemma E.1. Although the specified ranges for $d$ and $\alpha$ may appear somewhat restricted, they actually cover the cases commonly encountered in geostatistical applications, where Vecchia GPs are routinely employed. In such applications, the domain typically corresponds to a geological space with $d=1$ or $2$, and the regularity parameter $\alpha$ generally does not exceed $5/2$.
The main idea of the proof stems from the observation that, on an infinite grid, the geometric shapes of $\mathrm{pa}(X_{i})$ are identical. Therefore, the polynomial interpolation operators $P_{j},\forall j$ can be viewed as convolution operators, which can be further converted to pointwise multiplication via the Fourier transform. This approach greatly simplifies our analysis. However, we consider a finite grid, where the behavior near the boundary differs from that of an infinite grid. Polynomial interpolation on a finite grid exhibits “boundary effects”, which has to be treated separately. To address this, we use data augmentation. We expand the dataset $\mathcal{X}_{n}$ with artificially placed locations, so that any $X_{i}\in\mathcal{X}_{n}$ and its ancestors are sufficiently far from the boundary. Specifically, we first embed the finite grid $\mathcal{X}_{n}$ into a larger grid and then construct a layered norming DAG on the expanded grid, following a procedure analogous to Algorithm 1. We make this general strategy precise in the remainder of the section.
Our proof is divided into three parts. The first part describes how to extend a grid to mitigate boundary effects. The second part defines polynomial interpolation on an infinite grid and shows that it is equivalent to finite polynomial interpolation on the extended grid. In the final part, we derive upper bounds for the polynomial interpolation on infinite grid using Fourier analysis.
E.5.1 Extension of grid
We first introduce some notation. Without loss of generality, let the dataset $\mathcal{X}_{n}$ be a $n=(2^{r}+1)^{d}$ grid for some $r\in\mathbb{N}$ and choose $\gamma=2$, i.e.
@@ E.5.1 Extension of grid
Observe, that the above DAG structure satisfies Condition 1 with $c_{d}=1$ and Condition 2 with a constant $c_{N}>0$. Next, we augment each $\mathcal{N}_{j}$ with elements from $\mathcal{G}_{j}$ to form new sets $\tilde{\mathcal{N}}_{j}$, ensuring that none of the elements $x\in\mathcal{X}_{n}$ are affected by boundary effects. Specifically, for all $X_{i}\in\mathcal{X}_{n}$, let $\mathcal{A}(X_{i})$ be the collection of all ancestors of $X_{i}$, i.e., the collection of all elements of $\mathcal{X}_{n}$ such that there exists a directed path from that element to $X_{i}$. Then for all $X\in\mathcal{A}(X_{i})\cap\mathcal{N}_{j}$, $j<\eta(i)$, under Condition 2 and recalling that $\gamma=2$, we have
$\|X_{i}-X\|_{\infty}\leq\sum_{j^{\prime}=j}^{\eta(i)}c_{L}\gamma^{-j^{\prime}}\leq 2c_{L}\cdot 2^{-j}.$ (E.20)
Next, we describe the iterative data augmentation procedure. Recall that in Algorithm 1 the parent sets for elements in layers $j\leq j_{0}$, with $j_{0}$ only depending on $\alpha$ and $d$, were initialized as $\mathrm{pa}(X_{i})=\cup_{j^{\prime}=0}^{j-1}\mathcal{N}_{j^{\prime}}$. In this case we do not use data augmentation and set $\tilde{\mathcal{N}}_{j}=\mathcal{N}_{j}$, $\forall\;,0\leq j\leq j_{0}$. For all $j_{0}<j\leq\eta(n)-1$, we iteratively augment each layer $\mathcal{N}_{j}$ to $\tilde{\mathcal{N}}_{j}$ as
@@
is defined in the same way as in Algorithm 1 and satisfies Condition 2. This also means that relation (E.20) holds for all $X_{i}\in\tilde{\mathcal{X}}$ as well.
For all $X_{i}\in\tilde{\mathcal{X}}_{n}$ and $j_{0}\leq j\leq\eta(i)-1$, the new layer $\tilde{\mathcal{N}}_{j}$ includes all grid points in $\mathcal{G}_{j}$ within $4c_{L}\cdot 2^{-j}-4c_{L}\cdot 2^{-\eta(i)}\geq 2c_{L}\cdot 2^{-j}$ distance of $X_{i}$. In view of (E.20), all potential ancestors of $X_{i}$ in the $j$th layer lie within a distance of $2c_{L}\cdot 2^{-j}$ from $X_{i}$, hence the DAG structure of Algorithm 4 eliminates all boundary effects for $X_{i}\in\cup_{j=j_{0}}^{\eta(n)}\tilde{\mathcal{N}}_{j}\subset\tilde{\mathcal{X}}_{n}$.
1Input grid data $\mathcal{X}_{n}$ with $n=\tilde{n}^{d}$ and order $l$. Set $m={l+d\choose l}$ and let $r$ be the largest integer such that $\tilde{n}\geq 2^{r}+1$.
@@
10   Augment $\mathcal{N}_{j}$ to $\tilde{\mathcal{N}}_{j}$ according to (E.21).
@@
20         Order the projection of $\cup_{0\leq j^{\prime}\leq j-1}\tilde{\mathcal{N}}_{j^{\prime}}$ onto the $h$th coordinate by increasing distance to $X[h]$, denoting the resulting sequence as $B_{h}=(x_{s(1),h},x_{s(2),h},\cdots,x_{s(n^{\prime}),h})$.
@@
We note that, instead of data augmentation, one could also trim the dataset to eliminate boundary effects, sacrificing only a negligible fraction of the data (vanishing as $n\rightarrow\infty$). This can be done in a manner analogous to data augmentation and is left as an exercise for the reader.
E.5.2 Polynomial interpolation on infinite grid
With boundary effects eliminated via data augmentation, we can define the polynomial interpolation operator on the infinite grid and align it with the finite-grid operators $P_{j},j\in\mathbb{N}$.
@@ E.5.2 Polynomial interpolation on infinite grid
$\mathbb{K}=\{k=(k[1],k[2],\cdots,k[d)]^{T}\in\mathbb{Z}^{d}:|k|\geq 1,0\leq k_{j}\leq 1\}.$ (E.22)
@@
Furthermore, let us denote by $\mathrm{pa}(k/2)$ the corner parent set defined in (16) for $k/2\in\mathcal{N}_{j}$ and $\cup_{0\leq j^{\prime}\leq j-1}\mathcal{N}_{j^{\prime}}$ a sufficiently large subset of $\mathbb{Z}^{d}$. Recalling the definition of the Vandermonde matrix $V_{A}$ from (B.1), we define a linear operator $\mathrm{In}_{k}$ ($\mathrm{In}$” taking the name from interpolation) on functions $f:\mathbb{Z}^{d}\rightarrow\mathbb{R}$ as
$\mathrm{In}_{k}(f)(z)=v_{k/2}^{T}V_{\mathrm{pa}(k/2)}^{-1}\bm{f}(z+\mathrm{pa}(k/2)),\qquad\forall z\in\mathbb{Z}^{d},$ (E.23)
@@
$P(f)(z)=\Bigg\{\begin{aligned} &f(z/2),&&\text{if}\,z[j]\text{ is even},\;\forall 1\leq j\leq d,\\ &\mathrm{In}_{k}(f)(\lfloor z/2\rfloor),\;&&\text{else, with $k=z-2\lfloor z/2\rfloor$, }\end{aligned}$ (E.24)
@@
This part aligns an arbitrarily function $f$ defined on the finite grid $\tilde{\mathcal{X}}_{n}$ with an appropriate function $f^{\prime}$ defined on the infinite grid $\mathbb{Z}^{d}$, such that, under proper scaling, both functions yield the same polynomial interpolation over the grid.
Specifically, for all $0\leq j_{1}\leq j_{2}\leq\eta(n)-1$ and $f\in L_{(}\tilde{\mathcal{N}}_{j_{1}})$, the objective of this paragraph is to find an integer $j_{3}\in\mathbb{N}$ satisfying $j_{1}\leq j_{3}\leq j_{2}$ and a function $f^{\prime}\in L_{1}(\mathbb{Z}^{d})$, such that
$P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)(x)=P^{j_{2}-j_{3}}(f^{\prime})(2^{j_{2}+1}x),\;\;\forall x\in\tilde{\mathcal{X}}_{n}\cap\mathcal{G}_{j_{2}+1},$ (E.25)
@@
$\|f^{\prime}\|_{\infty}\lesssim\|f\|_{\infty}.$ (E.26)
@@
Let $j_{3}=j_{1}$, then (E.26) is immediate. Furthermore, in view of Algorithm 4, for all $X_{i}\in\cup_{j=j_{1}}^{\eta(n)}\tilde{\mathcal{N}}_{j}\subset\cup_{j=j_{0}}^{\eta(n)}\tilde{\mathcal{N}}_{j}$, its ancestors are in $\tilde{\mathcal{X}}_{n}$. Therefore, $X_{i}$ has no boundary effect, implying directly (E.25).
For $j_{1}<j_{0}$, let us define an intermediate function $f_{j_{0}}\in L_{1}(\cup_{j=0}^{j_{0}}\tilde{\mathcal{N}}_{j})$ as
@@
Let $j_{3}=j_{0}$, then $\|f^{\prime}\|_{\infty}=\|f_{j_{0}}\|_{\infty}$ holds. Furthermore, since $j_{0}$ only depends on $d,\alpha$ and $P_{j},\forall j$ has bounded $L_{\infty}$ norm, we also have $\|f_{j_{0}}\|_{\infty}\lesssim\|f\|_{\infty}$. Thus, the upper bound (E.26) is satisfied. Furthermore, since for arbitrary $X_{i}\in\cup_{j=j_{0}}^{\eta(n)}\tilde{\mathcal{N}}_{j}$, there is no boundary effect, equation (E.25) also holds.
@@
Assertions (E.25) and (E.26) align a function defined on the finite grid $\tilde{\mathcal{X}}_{n}$ with an appropriate function defined on the infinite grid $\mathbb{Z}^{d}$. The polynomial interpolation operator $P$ possesses a localization property in the sense that the value of $P(f)(x)$ depends only on $f(x^{\prime})$ for a limited number of points $x^{\prime}$ that are near $x$. Building on this idea, by introducing a surrogate function on $\mathbb{Z}^{d}$, we can derive sharp bounds for the operator norms, which are suitable analysis with Fourier methods. Details are provided below.
For $j_{3}$ and $f^{\prime}\in L_{1}(\mathbb{Z}^{d})$ satisfying (E.25) and (E.26), and $X_{i}\in\tilde{\mathcal{X}}_{n}$, define the function $f_{i}^{\prime}\in L_{1}(\mathbb{Z}^{d})$ as
@@
In view of assertion (E.20) , the set $\{x:\|x-X_{i}\|_{\infty}\leq 2c_{L}\cdot 2^{-j_{3}}\}$ contains all the ancestors of $X_{i}$ that belong to the layers $\cup_{j=j_{3}}^{j_{2}-1}\tilde{\mathcal{N}}_{j}$.
@@
$P_{j_{2}}P_{j_{2}-1}\cdots P_{j_{1}}(f)(X_{i})=P^{j_{2}-j_{3}}(f^{\prime})(2^{j_{2}+1}X_{i})=P^{j_{2}-j_{3}}(f_{i}^{\prime})(2^{j_{2}+1}X_{i}),$ (E.27)